can you write an equation for each line wit the given point and slope. express the squationin slope intercept form
(2,5) m=5 and (-1/2,0)m=3

Answer by stanbon(75887)

can you write an equation for each line wit the given point and slope. express the squationin slope intercept form
(2,5) m=5
Form: y = mx + b
You must determine "m" and "b"
m = 5
The point tell you that y = 5 when x = 2, so you can solve for "b":
5 = 5*2 + b
b = -5
Equation:
y = 5x - 5
-----------------------------------------
and (-1/2,0)m=3
Form: y = mx + b
0 = 3(-1/2) + b
0 = (-3/2) + b
b = (3/2)
-------------------
Equation:
y = 3x + (3/2)
=====================
Cheers,
Stan H.